OpenStudy (anonymous):

What is the equation of the line in slope–intercept form that passes through the point (0, 3) and is perpendicular to the line y = 2x – 6?

OpenStudy (anonymous):

y = 2x + 6 has slope 2, therefore any line perpendicular to it will have slope -1/2 (the inverse reciprocal). The point (0,3) lies on the y-axis and is therefore the desired line's y-intercept. Since you need the equation in slope-intercept form, you've just got to throw the numbers into the equation. y = mx + b m = -1/2 b = 3 Therefore, DRUM ROLL PLEASE, y = (-1/2)x + 3 Hope this helped!

OpenStudy (anonymous):

the line whose eqn we need to find is perpendicular to line y = 2x – 6 slope of line y = 2x – 6 is m=2 so slope of required line is negative inverse of given line i.e. -1/2 also the line passes through point (0,3) eqn of line with given and a point is y-y1=m(x-x1) where m=-1/2 and ( x1,y1)=(0,3) plug in these values y-y1=m(x-x1) y-3=-1/2(x-0) y-3=-x/2 y=(-x/2) + 3 is the required eqn of line